OpenStudy (clobber02):

Which is the side length of a cube with a volume of 64 m3?

OpenStudy (clobber02):

plz help me

OpenStudy (clobber02):

@AlexandervonHumboldt2

OpenStudy (clobber02):

@pooja195 (Ambassador)

OpenStudy (clobber02):

I WILL MEDAL AND FAN

OpenStudy (clobber02):

@Chlobug

OpenStudy (clobber02):

@mathmate

OpenStudy (mathmate):

Side length = L m Volume = \(L^3~ m^3\) So solve for L in equation \(L^3 = 64 m^3\)

OpenStudy (clobber02):

thx

OpenStudy (clobber02):

What is the volume of this triangular prism?

OpenStudy (clobber02):

@mathmate

OpenStudy (clobber02):

?

OpenStudy (mathmate):

ALL prisms have the formula Volume = \(A_b*h\) where \(A_b = \) area of base (of prism) h = height of prism Area of a triangle = \(\dfrac {base \times height}{2}\) where base = length of base of triangle height = height of triangle

OpenStudy (clobber02):

hhmm

OpenStudy (clobber02):

i dont get it

OpenStudy (clobber02):

bxh /2 ?

OpenStudy (clobber02):

B \[B \times H \div 2\]

OpenStudy (clobber02):

@mathmate

OpenStudy (clobber02):

is the answer 12? @mathmate

OpenStudy (mathmate):

Yes, BH/2 for the volume of a triangle (or base of the prism)

OpenStudy (clobber02):

thank you so much you are really helping me!!!

OpenStudy (clobber02):

wait 12 is not an answer

OpenStudy (clobber02):

A. 108 cm3 B. 54 cm3 C. 216 cm3 D. 532 cm3

OpenStudy (mathmate):

Yes, the area of the base of the prism is 12. Do not forget, question wants the \(volume\) of the prism. Use the other formula to finish the problem.

OpenStudy (mathmate):

Sorry, I am not interested in the answer choices. I am interested what you calculate and how you got it.

OpenStudy (clobber02):

what?

OpenStudy (clobber02):

what are u talking about

OpenStudy (mathmate):

Sorry, I am not interested in the answer choices. I am interested in what you calculate and how you got it.

OpenStudy (clobber02):

i hate u

OpenStudy (mathmate):

I am very sorry, guess it hard to please everyone. Have a good evening anyway!

OpenStudy (clobber02):

plz i really need your help